Find u from the differential equation and initial condition.



du/dt= e^(1.5t-1.3u), u=0 1.3



Find u=?




1
Expert's answer
2022-03-13T18:49:43-0400

"\frac{du}{dt} = e^{1.5t-1.3u}" , "u(0) = 1.3"


"=>\frac{du}{dt}=e^{1.5t}e^{-1.3u}"


"=>e^{1.3u}du=e^{1.5t}dt"


Interacting both sides, we have;


"=>\frac{e^{1.3u}}{1.3}=\frac{e^{1.5t}}{1.5} + c"


Recall that "u(0)=1.3"


"=>\frac{e^{1.3(1.3)}}{1.3}=\frac{e^{1.5(0)}}{1.5} + c"


"=>\frac{e^{1.69}}{1.3}=\frac{e^{0}}{1.5} + c"


"=>4.17=0.67+c"


"=>c=4.17-0.67=3.5"


"=>\frac{e^{1.3u}}{1.3}=\frac{e^{1.5t}}{1.5}+3.5"


Multiply through by 1.3. we have;


"=>e^{1.3u}=\frac{13e^{1.5t}}{15}+\frac{91}{20}"


"=>e^{1.3u}=\frac{52e^{1.5t}+273}{60}"


Take natural log of both sides


"=>lne^{1.3u}=ln(\frac{52e^{1.5t}+273}{60})"


"=>1.3u=ln(\frac{52e^{1.5t}+273}{60})"


"=>u=\frac{10ln(\frac{52e^{1.5t}+273}{60})}{13}"